What is the effective half life of Tc-99m.? & If the biological half life of Tc-99m is 10h and the physical half life is 6 h.
1
Expert's answer
2011-06-10T06:18:30-0400
We can calculate the effective half-life using formula
1/te = 1/tb + 1/tph
Hence te=10*6/(10+6)=3.75 h